OpenStudy (anonymous):

I just took a test, I got a 28 out of 40, can someone tell if I passed? lol

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

(28/40)*100 = 0.7*100 = 70% so you got a C-

OpenStudy (anonymous):

70%

OpenStudy (anonymous):

awesome,I did better then I thought haha, thanks guys.

jimthompson5910 (jim_thompson5910):

For the most part, I would say you passed, but I'm not sure what your teacher considers passing or not (most would say 60% or more is passing, but some might have the bar higher)